Hi! Can you help me with this problem?
One number is 4 times another number. The larger number is also 87 more than the smaller number. Find the numbers.
If you could help me it would be very much appreciated.
Thanks.

Found 2 solutions by fractalier, stanbon:
Answer by fractalier(6550)

Let x be the smaller and y be the larger.
so we have
y = 4x and
y - x = 87
Now substituting we have
4x - x = 87
3x = 87
x = 29
y = 4x = 116

Answer by stanbon(75887)

One number is 4 times another number. The larger number is also 87 more than the smaller number. Find the numbers.
Let "s" be the smaller and "l" be the larger number.
EQUATIONS:
l=4s
s=l-87
Substitute to get:
s=4s-87
-3s=-87
s=29
l=29+87=116
Cheers,
Stan H.